One leg of a right triangle is 7cm shorter than the other leg. How long should be the shorter leg to ensure the hypotenuse is at least 13?
Answer by Cromlix(4381)

Hi there,
Make 'other leg' = x
One leg = x - 7
Using Pythagoras' Theorem.
x^2 + (x - 7)^2 = 13^2
x^2 + x^2 - 14x + 49 = 169
Rearrange and collect like terms.
2x^2 - 14x - 120 = 0
Divide by 2 throughout (for ease)
2(x^2 - 7x - 60) = 0
2(x + 5)(x - 12) = 0
x + 5 = 0
x = -5 (Disregard as -ve)
x - 12 = 0
x = 12
Other leg = 12 cm
One leg = 5 cm
Shorter leg = 5 cm
Hope this helps :-)